OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 2x+3y=5 and contains the point (-2,1).

OpenStudy (campbell_st):

rewrite you given line in point slope form \[y =\frac{ -2x}{3} + 5\] this line has a gradient of -2/3 so the equation of a parallel line will be \[y = \frac{-2x}{3} + b\] substitute x = -2 and y = 1 to find the value of b

OpenStudy (anonymous):

a. 2x-3y =7 b. 3x-2y=7 c. 3x+2y=5 d. 2x+3y=-1

OpenStudy (anonymous):

d?